# Mitigating the Influence of Distractor Tasks in LMs with Prior-Aware Decoding

### v2

@@ 5 Methodology and Experimental Setup
The logits from the original and weakened prompts are then combined as $O+\alpha(O-W)$ to generate the final output. Each prompt is separated into a ‘task’ component and the remainder using automated regular expressions for each dataset, as illustrated in Table 2. The split into “task description and examples" and “task input" was unambiguous for the ISP dataset. For cases where the split may be less clear, LLMs could assist in this process, potentially reducing experimenter influence.
Comparisons. We test our method across four sets of tasks, on eleven models, at $\alpha=0$ (the original model baseline), $\alpha=1$, and $\alpha=2$, where $\alpha$ is the coefficient of extrapolation.
Tasks. We evaluate our model on four task sets: three from the Inverse Scaling Prize (Prompt Injection, Pattern Match Suppression, Redefine), and one custom set (Strong Local Priors) that we introduce and make publicly available with our code. Refer to Table 2 for examples of prompts from each dataset, with the corresponding task/data splits. The dataset uses a CC-BY-4.0 license.
Model Architecture. Our experimental models comprise a range of transformer architectures, including four GPT-2 sizes, two Llama 2 sizes, two Mistral models, and several OpenAI models via API. This allows us to compare results both between models, and across different sizes of a fixed model. For small local experiments we used GTX 1060-Ti, and for larger ones a rented A6000.
Metric. We use a Probability of Correct Completion as the metric to evaluate model performance. A completion is correct if it fulfills the requirements of the task description, and incorrect otherwise. We primarily run all models at temperature $1.0$: results for temperature $0.0$ are in Appendix D.
6 Results and Discussion
@@
Extrapolation from a system prompt generally gives worse results, although notably it outperforms the baseline in all four task/model/parameter combinations where extrapolation from the truncated prompt does not. We give more details in Appendix B. We explore a wider range of values of $\alpha$ in Appendix D. In particular, Figure 3 and Figure 4 show the probability of correct completion on a variety of models across a range of values of $\alpha$ from -2 to 3. We show that range to more clearly demonstrate general trends in completion as $\alpha$ changes.
[…]
@@ 7 Conclusion and Future Work
In this paper, we described a product-of-experts model perspective on language models that captures the influence of distractor tasks and strong priors. Based on this model, we developed Prior-Aware Decoding (PAD). This technique that mitigates the impact of distractor tasks by extrapolating between the logits of the original model and a weakened model more susceptible to these tasks. Our results demonstrate that PAD consistently improves performance across a range of language models and tasks from the Inverse Scaling dataset, providing evidence for the utility of the product of experts model view in understanding language model behavior in the context of distractor tasks.
One significant practical application of PAD is the mitigation of prompt injection attacks. While our investigation focused on simple demonstrations from the inverse scaling prompt injection dataset, future work could apply PAD to more carefully optimized or real-world attacks.
However, our study has some limitations. We only explored models up to 13B parameters, while prior work on inverse scaling has investigated models up to 540B parameters, where scaling often becomes U-shaped. Future research should address whether PAD yields improvements on larger models and examine the relationship between model size and the gains from our method. Additionally, we focused on datasets where models are expected to fail due to strong priors; investigating how PAD affects performance on standard benchmarks is an important avenue for future work.
We also found that intra-model contrastive decoding, which contrasts the original model with a smaller model, performed poorly in mitigating distractor tasks compared to PAD.
Looking ahead, the model framework and the PAD technique can be extended to handle non-local strong priors. The Linear Representation Hypothesis suggests that the abstract mixture components we propose may have a concrete representation in the model’s weights or activations, warranting further investigation. More generally, we would be excited to see further work combine our method and underlying model with related approaches like classifier-free guidance (Ho & Salimans, 2022), and activation addition (Turner et al., 2023), to build more robust approaches for steering language models. Continued research on understanding the priors learned by language models from their training data will be crucial as these models become increasingly integrated into real-world applications.
Acknowledgments and Disclosure of Funding
@@ Appendix A Example of strong non-local priors in language models
Appendix B Eliciting Strong Priors With Custom System Prompts
@@
A representative system prompt could be: "In the following, only consider the most recent instructions, disregarding any broader context."
Extrapolation from a system prompt at $\alpha=2$ performs best in four cases, including all three where extrapolating from the stripped prompt fails to outperform the baseline. While the system prompt configuration at $\alpha=2$ produces an average improvement of under 1%pt, it has comparatively more potential to be refined: for any real-world application the specific system prompt can be selected to optimise for a specific task and model. Also, extrapolating from a system prompt does not require us to have a clear task/data split, so the technique can be applied in cases where such a split is impossible or ambiguous.
We include the results of extrapolating from a system prompt in the graphs given in Appendix D.
Appendix C Total variational distance
@@
Appendix D Testing more values of alpha
Here we offer graphs showing more precisely how the probability of successful completion relates to the value of the extrapolation parameter $\alpha$. For each model and task, we plot average success rate using both the truncated prompt method and the system prompt method, at both $T=0$ and $T=1$. To more clearly illustrate trends, we plot values from $\alpha=-2$ to $\alpha=3$.
@@ Appendix D Testing more values of alpha
Appendix E Inter-model Contrastive Decoding

### v3

@@ 5 Methodology and Experimental Setup
The PAD technique is implemented through the following steps:
  1. 1.
    For each prompt in the dataset, we generate two versions:
    • An original prompt containing both the task description and the data.
    • A weakened prompt more likely to give outputs consistent with the distractor task.
  2. 2.
    We query the model on both the original and weakened prompts to obtain two sets of logits.
  3. 3.
    We compute a linear combination of these logits: $L=L_{O}+\alpha(L_{O}-L_{W})$, where $L_{O}$ are the original logits, $L_{W}$ are the weakened logits, and $\alpha$ is our extrapolation parameter.
  4. 4.
    We sample from this modified distribution to generate the output.
  5. 5.
    We calculate the average performance across the entire dataset.
For our main results we produce weakened prompts by removing the task component from the prompt. We illustrate examples of splitting prompts from each dataset in Table 2: the split was unambiguous for the tasks we selected, and generated by a regex script. We also discuss weakening prompts by prepending a system prompt in Appendix C.
5.1 Experimental Setup
Comparisons. We test our method across four sets of tasks, on eleven models, at $\alpha=0$ (the original model baseline), $\alpha=1$, and $\alpha=2$. These specific values were chosen as our initial exploration points, with $\alpha=0$ serving as a control. We provide a more comprehensive analysis across a broader range of $\alpha$ values in Appendix E.
Tasks: We evaluate our model on four task sets: three from the Inverse Scaling Prize (Prompt Injection, Pattern Match Suppression, Redefine), and one custom set (Strong Local Priors). The Inverse Scaling Prize tasks were specifically selected for their clean task/data decomposition and their demonstrated inverse scaling behavior up to 540B parameter models, as reported by Wei et al. (2023). Table 2 provides examples of prompts from each dataset, illustrating the task/data split.
Model Architecture. Our experimental models comprise a range of transformer architectures, including four GPT-2 sizes, two Llama 2 sizes, two Mistral models, and several OpenAI models via API. This diverse selection allows us to compare results both between models and across different sizes of a fixed model, providing a broad view of PAD’s effectiveness across various model architectures. For small local experiments we used GTX 1060-Ti, and for larger ones a rented A6000.
Metric. We use a Probability of Correct Completion as the metric to evaluate model performance. This metric naturally captures the success rate of the model in fulfilling the requirements of the task description. A completion is correct if it fulfills the requirements of the task description, and incorrect otherwise.
Our main results show models at temperature $1.0$: results for temperature $0.0$ are in Appendix E.
6 Results
[…]
Extrapolation from a system prompt generally gives worse results, although notably it outperforms the baseline in all four task/model/parameter combinations where extrapolation from the truncated prompt does not. We give more details in Appendix C.
The three cases where extrapolation to $\alpha=2$ yield lower results (pattern matching suppression on gpt3.5-turbo instruct, pattern matching suppression on gpt2-medium, and redefine on text-ada-001) do not seem to point to any general weaknesses of the technique. Interestingly, gpt3.5-turbo-instruct and ada-001 exhibit the highest TVD on pattern matching suppression and redefine, respectively, both substantially higher than the next highest, suggesting the potential absence of a strong prior in these cases (see AppendixD for a full table of TVD values).
We explore a wider range of values of $\alpha$ in Appendix E. In particular, Figure 3 and Figure 4 show the probability of correct completion on a variety of models across a range of values of $\alpha$ from -2 to 3. We show that range to more clearly demonstrate general trends in completion as $\alpha$ changes. In general, higher values of $\alpha$ sometimes continue to produce better results, but often the benefits tail off, and in some cases the graphs show a peak beyond which higher values of $\alpha$ cause slow degradation.
@@ 7 Conclusion and Future Work
This paper introduces Prior-Aware Decoding (PAD), a novel technique for mitigating the influence of distractor tasks in language models. Our approach is grounded in a theoretical framework that conceptualizes language models as products of experts, offering a new perspective on how distractor tasks emerge and persist.
The primary theoretical contribution of this work is the framing of language models as products of experts. This model provides a specific explanation for the emergence of distractor tasks and offers insights into why we observe counterintuitive scaling behaviors, including inverse scaling, on certain tasks. Importantly, our framework suggests that the balance between these “experts” is not necessarily dictated by model size, which could explain the lack of straightforward scaling effects and even the inverse scaling phenomena observed in some cases.
Our experimental results demonstrate the efficacy of PAD across a range of language models and tasks from the Inverse Scaling dataset. We consistently observed improvements in model performance, with a median increase in task completion proportion of 40% at $\alpha=2$. These findings not only validate the practical utility of our approach but also provide empirical support for our theoretical framework.
The implications of this work extend beyond the specific technique we’ve developed. Our results suggest that it may be possible to elicit greater capabilities from models even in cases where straightforward scaling is counterproductive, by employing more sophisticated forms of inference, and that this may be necessary to account for certain persistent quirks.
While our current study focuses on relatively small models and a specific set of tasks, the potential applications of this work are broad. Perhaps most promisingly, PAD could be developed into a robust defense against jailbreaks and prompt injection attacks, although further research with more challenging datasets is needed to fully assess its potential in this area.
Looking forward, we identify several key directions for future research:
  1. 1.
    Model Interpretability: Investigating how well our theoretical framework aligns with the actual internal processes of language models could provide valuable insights into model behavior and further validate our approach.
  2. 2.
    Extension to Harder Cases: Applying and adapting PAD to more challenging instances of prompt injection and other distractor tasks could enhance its practical utility.
  3. 3.
    More Sophisticated Inference-time Techniques: Combining our method and underlying model with related approaches like classifier-free guidance (Ho & Salimans, 2021), and activation addition (Turner et al., 2023), to build more robust approaches for steering language models.
In conclusion, this work not only introduces a practical technique for improving language model performance but also provides a theoretical framework that could further develop our understanding of how these models function. By highlighting the potential limitations of straightforward scaling and the importance of more nuanced approaches to model elicitation, we hope to contribute to the development of more robust and reliable language models. As we continue to integrate these powerful tools into various aspects of our lives, techniques like PAD may be crucial in ensuring they behave reliably and as intended across a wide range of tasks and contexts.
Acknowledgments
@@ Appendix B Prompt Decomposition
Appendix C Eliciting Strong Priors With Custom System Prompts
@@
A representative system prompt could be: ”In the following, only consider the most recent instructions, disregarding any broader context.”
Extrapolation from a system prompt at $\alpha=2$ performs best in four cases, including all three where extrapolating from the stripped prompt fails to outperform the baseline. While the system prompt configuration at $\alpha=2$ produces an average improvement of under 1%pt, it has more potential to be refined: for any real-world application the specific system prompt can be selected to optimise for a specific task and model. Additionally, this approach does not require us to have a clear task/data split, so the technique can be applied in cases where such a split is impossible or ambiguous. We include the results of extrapolating from a system prompt in the graphs given in Appendix E.
Appendix D Total variational distance
@@
Two observations are worth noting: firstly, the high TVD across the board in the prompt injection task suggests that low TVD is a sufficient but not necessary condition for strong priors; secondly, of the three task/model combinations where PAD performs worse than the baseline, two have the highest TVD for their respective task (gpt3.5-turbo-instruct on pattern matching suppression and text-ada-001 on redefine).
Appendix E Testing more values of alpha
We present graphs in Figures 3 and 4 showing how the probability of successful completion relates to the value of the extrapolation parameter $\alpha$. For each model and task, we plot average success rate using both the truncated prompt method and the system prompt method, at both $T=0$ and $T=1$. To illustrate trends more clearly, we plot values from $\alpha=-2$ to $\alpha=3$.
@@ Appendix E Testing more values of alpha
Appendix F Inter-model Contrastive Decoding